# How does Alignment Enhance LLMs’ Multilingual Capabilities? A Language Neurons Perspective

### v1

@@ 4.3 Layer-wise Functionality Analysis
    Multilingual Output Space Transformation: The model transfers features into the multilingual output space in this stage in preparation for generating the final output. In this part, the number of language neurons reaches a peak again, while the number of language-agnostic neurons drops to the lowest point.
@@
    Vocabulary Space Outputting: In the last layer, the model maps vectors of different languages into a shared vocabulary space to generate outputs. The number of both language-related and language-agnostic neurons rises sharply, while the language-specific neurons are fewer than those in several previous layers.
Meanwhile, the distribution of different types of neurons aligns with the conclusions from existing studies mentioned above. Overall, we can find that the number of neurons varies correspondingly with the different inference stages of LLMs.
4.4 Layer-wise Neuron Changes Analysis
In §4.3, we investigate the fundamental distribution of neurons and the basic partitioning of the model. We further analyze the changes in different types of neurons before and after multilingual alignment. Based on the four functional stages in LLMs, we quantify the layer-wise changes ($\Delta$) in the number of different types of neurons. Figure 4 presents the results for language-specific neurons, language-related neurons, and language-agnostic neurons.
@@ 4.4 Layer-wise Neuron Changes Analysis
During the multilingual understanding stage, the number of language neurons increases, while language-agnostic neurons decrease. In the subsequent shared semantic space reasoning stage, language-agnostic neurons increase substantially, whereas language neurons remain stable and nearly absent. Then, in the third stage, as language-agnostic neurons decrease, language neurons increase overall. Additionally, we notice that the number of language-related neurons show an upward trend. Finally, in the last stage, the number of language-agnostic increases significantly in the aligned model, accompanied by a reduction in language neurons. We also report the results of different checkpoints during the alignment process in Appendix F.
@@
Overall, we find that language neurons and language-agnostic neurons exhibit generally opposite trends across different layers, which corresponds to the characteristics of LLMs at different stages of inference. Especially, at the last stage, language-agnostic neurons play an more important role than language neurons. Multilingual alignment facilitates more effective activation of the appropriate neurons at each stage, thereby improving the model’s capability to handle multilingual tasks.
4.5 Macroscopic Analysis of Different Types of Neurons
@@
We further conduct macroscopic analysis for different types of neurons. In our neuron identification algorithm introduced in §3.3, the number of languages that share a specific neuron is an attribution characterizing all types of activated neurons. Since our study involves 10 languages, the valid range of $N$ is from $1$ to $10$. Among these, values of $N$ from $2$ to $9$ correspond to language-related neurons. As special cases in our work, $N=1$ represents language-specific neurons, while $N=10$ corresponds to language-agnostic neurons.
@@
Figure 5: Changes in the number of neurons shared by $N$ languages after alignment.
@@
We report the changes in the number of neurons after multilingual alignment for each value of $N$ ranging from 1 to 10 in Figure 5. The results show a decrease in the number of language-specific neurons, while an increase in the number of language-related neurons, which are shared across multiple languages. This indicates that multilingual alignment encourages LLMs to develop and utilize more shared language-related neurons, rather than on language-specific neurons, which are applicable to only a single language. Meanwhile, during the alignment process, the model improves its understanding of task-relevant common knowledge. Therefore, the overall number of language-agnostic neurons also increases significantly.
Table 1: Accuracy of the MistralMathOctopus base model and aligned model on MGSM. "X/Y $\Rightarrow$ T" indicates that languages X and Y are used for multilingual alignment.
@@
4.6 Spontaneous Multilingual Alignment Analysis
The important “Spontaneous Multilingual Alignment” phenomenon is first revealed and discussed by Zhang et al. (2024), which means that conducting alignment in a small number of languages significantly improves multilingual alignment even between English and many languages unseen during the alignment process. We further analyze this phenomenon in our experiments. As shown in Table 1, spontaneous multilingual alignment also emerges under the multilingual alignment strategy employed in our study. Except for the languages used for alignment, LLMs exhibit notable performance gain in other unaligned languages.
To understand how multilingual alignment generalizes to other languages, we analyze the changes in different types of neurons before and after multilingual alignment based on our method.
@@ 4.6 Spontaneous Multilingual Alignment Analysis
Taking the case of “zh/de $\Rightarrow$ en” as a representative example, we report the average results in Table 4.6. For the trained languages, the number of language-specific neurons decreases, while the number of language-related neurons increases. This indicates that the aligned languages tend to utilize more language-related neurons shared with other languages rather than exclusive language-specific neurons. Moreover, we extend this analysis to languages other than the trained languages and observe a similar phenomenon. These findings indicate that multilingual alignment facilitates the use of language-related neurons while reducing the reliance on language-specific neurons in both trained and other unseen languages. We hypothesize that the new language-related neurons shared with trained languages contribute to the performance improvement on other unseen languages.
@@
4.7 Further Analysis
Uniqueness of English
Since current LLMs are primarily pretrained on English data, English is often regarded as playing a special role within LLMs (Wendler et al., 2024). In our experiments, we also observe that English exhibits markedly different characteristics compared to other non-English languages. Based on the identified neurons in our work, in Figure 2, we can find that deactivating the language neurons of English has a negligible impact on the model’s performance in English, which is entirely different from the behavior observed in other languages. This is also consistent with the results of Tang et al. (2024). Furthermore, based on this finding, we quantify the number of language neurons for English and non-English languages based on the MistralMathOctopus base model (Table 4.6). Our analysis reveals that English has significantly fewer neurons than other languages, both in terms of language-specific and language-related neurons. We hypothesize that this is due to the fact that English actually possesses numerous language-related neurons. And since English serves as a pivot language, these language-related neurons are likely shared with almost all other languages. It causes them to be confounded with language-agnostic neurons.
@@ 4.7 Further Analysis
We discuss the stability of neuron distributions across different data domains, as well as before and after alignment. To quantify the stability of neuron distributions, we compute the neuron overlap ratio in both settings, with the results summarized in Table 4.6. We can find that although the exact positions of a few language neurons may vary across different settings, the positional distribution of most language neurons remains stable. This also indicates good reliability and generalization of the language neurons identified under fixed hyperparameters.
5 Conclusion
In this work, we systematically investigate the multilingual alignment from the perspective of language neurons. We propose a new language neuron identification algorithm based on entropy and probability value, which detects the language-specific neurons, language-related neurons, and language-agnostic neurons in LLMs. The validity of the identified neurons is confirmed through deactivation ablation experiments. Furthermore, we examine the multilingual alignment mechanism by analyzing the roles of different types of neurons. Based on their distributional characteristics, we categorize LLMs’ internal process into four functional parts. Our analysis reveals that multilingual alignment enhances the model’s utilization of the corresponding types of neurons across different functional parts. Meanwhile, we find that alignment promotes a greater reliance on shared language-related neurons across languages, rather than on language-specific neurons. We also explore the phenomenon of spontaneous multilingual alignment. Additionally, we provide further analysis and more empirical results based on the preceding findings.
Appendix A Limitations
Appendix B Implementation Details
Our experiments are conducted on 4 NVIDIA RTX A6000 GPUs or 4 NVIDIA GeForce RTX 3090 GPUs. We use the TRL [von Werra et al., 2020] and DeepSpeed [Rasley et al., 2020] frameworks for preference alignment, and the vLLM engine Kwon et al. [2023] for inference.
B.1 MAPO
@@
B.2 Language Neurons Identification
@@
Appendix C Licenses for Used Assets
@@
Appendix D Deactivation Ablation Experiments
For the MistralMathOctopus model, Figure 6 presents the results of the deactivation ablation experiments conducted on MGSM using the neuron identification algorithm proposed by Tang et al. [2024]. Additionally, Figure 7 presents the results of the deactivation ablation experiments on the out-of-domain test set, MSVAMP. To verify the generalization of our identification algorithm, we also deploy the MetaMathOctopus to conduct the deactivation ablation experiment, with the result in Figure 8.
Appendix E Layer-wise distribution of the different types of neurons
In Figure 9(a) and 9(b), we separately display the layer-wise distribution of the different types of neurons of MistarlMathOctopus on MSVAMP and MetaMathOctopus on MGSM to validate the generalization of our finding discussed in §4.3.
Appendix F Layer-wise Changes in the Number of Different Types of Neurons During Alignment
[…]
Appendix G Spontaneous Multilingual Alignment
We report the accuracy of the base model on the MGSM and MSVAMP benchmarks in Table 5. The expected accuracy of random guessing is 50.0%. "X/Y $\Rightarrow$ T" indicates that languages X and Y are used for alignment. The best results for each language are highlighted. It is evident that models trained on multilingual translation data substantially outperform the original model across a wide range of languages, indicating that multilingual training significantly enhances the model’s multilingual capabilities.
@@ Appendix G Spontaneous Multilingual Alignment
Figure 6: PPL changes of MistralMathOctopus on MGSM when deploying the algorithm proposed by Tang et al. [2024].
@@
(c) Base - Language Neurons
(d) Aligned - Language Neurons
@@
(a) Base - Language Neurons on MGSM
(b) Aligned - Language Neurons on MGSM
Figure 8: PPL changes of MetaMathOctopus on MGSM after deactivating language neurons. “Base” indicates the results of the base model. “Aligned” indicates the results of the aligned model.
@@
Figure 13: Layer-wise changes in the number of language-agnostic neurons of MistralMathOctopus during the alignment.

### v2

@@ Introduction
    Multilingual Output Space Transformation: The model transfers features into the multilingual output space in this stage in preparation for generating the final output. In this part, the number of both language-specific and language-related neurons reaches a peak again, while the number of general neurons drops to the lowest point.
@@
    Vocabulary Space Outputting: In the last layer, the model maps vectors of different languages into a shared vocabulary space to generate outputs. The total number of language-specific neurons and language-related neurons increases, but there is an opposite trend between language-specific and language-related neurons. Additionally, unlike assumptions in previous work (Wendler et al. 2024; Zhao et al. 2024b), the number of general neurons reaches maxima at this stage, consistent with the behavior in the early layers. We therefore speculate that this may be related to the shared vocabulary across different languages.
Meanwhile, the distribution of different types of neurons aligns with the conclusions from the existing studies mentioned above. Overall, the number of neurons varies correspondingly with the different inference stages of LLMs.
Layer-wise Neuron Changes Analysis
We further analyze the changes in different types of neurons before and after alignment. Based on the four functional stages above, we quantify the layer-wise changes ($\Delta$) in the number of different types of neurons in Figure 4.
@@
During the multilingual understanding stage, the number of both language-specific and language-related neurons increases, while general neurons decrease. In the subsequent shared semantic space reasoning stage, general neurons increase substantially, whereas both language-specific neurons and language-related neurons remain stable and nearly absent.
In the third stage, as general neurons decrease, the sum of language-specific neurons and language-related neurons increase overall. Additionally, we notice that language-specific and language-related neurons show opposite trends. This reflects that the aligned model relies more on shared language-related neurons than language-specific neurons in the third stage, which is difficult to observe under the previous taxonomies.
Finally, in the last stage, the number of general neurons increases significantly in the aligned model, accompanied by a reduction in both language-specific and language-related neurons. This contradicts the conventional three-stage assumption, indicating that general neurons also play a significant role in the final vocabulary outputting stage. Additionally, combined with the observations in Figure 5, the model relies on a smaller set of language neurons with a higher degree of sharing in this stage. We also report the results of different checkpoints during the alignment process in appendix.
@@
Overall, we find the union set of language-specific and language-related neurons and general neurons exhibit generally opposite trends across layers, which corresponds to the characteristics of LLMs at different stages. Especially, at the last stage, general neurons play an more important role. Multilingual alignment facilitates more effective activation of the appropriate neurons at each stage, thereby improving the model’s capability to handle multilingual tasks.
Macroscopic Analysis of Neurons
@@
We further conduct macroscopic analysis for different types of neurons. In our neuron identification algorithm, the number of languages that share a specific neuron is an attribution characterizing all types of activated neurons. Since our study involves 10 languages, the valid range of $N$ is from $1$ to $10$. Among these, values of $N$ from $2$ to $9$ correspond to language-related neurons. As special cases in our work, $N=1$ represents language-specific neurons, while $N=10$ corresponds to general neurons.
@@
Figure 5: Changes in the number of neurons shared by NN languages after alignment.
@@
We report the changes in the number of neurons after multilingual alignment for each value of $N$ ranging from 1 to 10 in Figure 5. The results show a decrease in the number of language-specific neurons, while an increase in the number of language-related neurons, which are shared across multiple languages. This indicates that multilingual alignment encourages LLMs to develop and utilize more shared language-related neurons, rather than language-specific neurons, which are applicable to only a single language. Meanwhile, during the alignment, the model improves its task-relevant common knowledge. Therefore, the number of general neurons also increases significantly. Language-related neurons and language-specific neurons exhibit overall opposite trends, further highlighting the importance of a more fine-grained analysis of the two categories within our new taxonomy. In addition, we also report the overall neuron numbers for each value of $N$ before and after alignment in Figure 10.
@@
Table 1: Accuracy of the MistralMathOctopus base model and aligned model on MGSM. ”X/Y $\Rightarrow$ T” indicates that languages X and Y are used for multilingual alignment.
Spontaneous Multilingual Alignment Analysis
The “spontaneous multilingual alignment” phenomenon is first revealed and discussed by Zhang et al. (2024), which shows that conducting alignment in a small number of languages significantly improves multilingual alignment even between English and many languages unseen during the alignment process. We further analyze this phenomenon in our experiments. As shown in Table 1, spontaneous multilingual alignment also emerges under the multilingual alignment method employed in our study. Besides the languages used for alignment, LLMs exhibit notable performance gains in other unaligned languages. To understand how multilingual alignment generalizes to other languages, we analyze the changes in different types of neurons before and after multilingual alignment based on our method.
@@
Taking the case of “zh/de $\Rightarrow$ en” as a representative example, we report the average results in Table 2. For the trained languages, the number of language-specific neurons decreases, while the number of language-related neurons increases. This indicates the aligned languages tend to utilize more language-related neurons shared with other languages rather than exclusive language-specific neurons. Moreover, we extend this analysis to languages other than the trained languages and observe a similar phenomenon. These findings indicate multilingual alignment facilitates the use of language-related neurons while reducing the reliance on language-specific neurons in both trained and other unseen languages. We hypothesize that the new language-related neurons shared with trained languages contribute to the performance improvement on other unseen languages.
@@
Further Analysis
Uniqueness of English
Since current LLMs are primarily pretrained on English data, English is often regarded as playing a special role within LLMs (Wendler et al. 2024). In our work, we observe that English exhibits markedly different characteristics compared to other languages. Based on the identified neurons in our work, in Figure 2, deactivating the language neurons of English has a negligible impact on the model’s performance in English, which is entirely different from the behavior observed in other languages. This is consistent with the results of Tang et al. (2024). Furthermore, we quantify the sum of language-specific neurons and language-related neurons for English and non-English languages based on the MistralMathOctopus base model (Table 3).
Our analysis reveals that English has significantly fewer neurons than other languages, both in terms of language-specific and language-related neurons. We hypothesize that this is due to that English actually possesses numerous language-related neurons. And since English serves as a pivot language, these language-related neurons are likely shared with almost all other languages. It causes them to be also regarded as general neurons.
@@
We discuss the stability of neuron distributions across different data domains, as well as before and after alignment. To quantify the stability of neuron distributions, we compute the neuron overlap ratio in both settings, with the results summarized in Table 4. We can find that although the exact positions of a few neurons may vary across different settings, the positional distribution of most neurons remains stable. This also indicates good reliability and generalization of the language neurons identified under fixed hyperparameters.
Conclusion
In this work, we systematically investigate the multilingual alignment from the perspective of language neurons. First, based on the defects of the existing binary classification methodology, we propose a ternary classification methodology, which defines the language-specific neurons, language-related neurons, and general neurons. Then we propose a corresponding language neuron identification algorithm, which detects the above different types of neurons in LLMs.
Furthermore, we examine the multilingual alignment mechanism by analyzing the roles of different types of neurons. Based on their distributional characteristics, we categorize LLMs’ internal process into four functional parts. Our analysis reveals that multilingual alignment enhances the model’s utilization of the corresponding types of neurons across different functional parts. Meanwhile, we find that alignment promotes a greater reliance on shared language-related neurons across languages, rather than on language-specific neurons. This also corresponds to the phenomenon of spontaneous multilingual alignment.
Overall, we provide further analysis and valuable insights to better understand multilingual alignment and multilingual capabilities of LLMs.
Acknowledgement
We would like to thank the anonymous reviewers for their insightful comments. Shujian Huang is the corresponding author. This work is supported by National Science Foundation of China (No. 62376116, 62176120), research project of Nanjing University-China Mobile Joint Institute (NJ20250038), the Fundamental Research Funds for the Central Universities (No. 2024300507, 2025300390).
Appendix A Implementation Details
Our experiments are conducted on 4 NVIDIA RTX A6000 GPUs or 4 NVIDIA GeForce RTX 3090 GPUs. We use the TRL (von Werra et al. 2020) and DeepSpeed (Rasley et al. 2020) frameworks for preference alignment, and the vLLM engine (Kwon et al. 2023) for inference.
MAPO
@@ Appendix A Implementation Details
Language Neurons Identification
@@
Appendix B Licenses for Used Assets
@@
Appendix C Deactivation Ablation Experiments
For the MistralMathOctopus model, Figure 6 presents the results of the deactivation ablation experiments conducted on MGSM using the neuron identification algorithm proposed by Tang et al. (2024). Additionally, Figure 7 presents the results of the deactivation ablation experiments on the out-of-domain test set, MSVAMP. To verify the generalization of our identification algorithm, we also deploy the MetaMathOctopus to conduct the deactivation ablation experiment, with the result in Figure 8.
Appendix D Layer-wise distribution of the different types of neurons
In Figure 9(a) and 9(b), we separately display the layer-wise distribution of the different types of neurons of MistarlMathOctopus on MSVAMP and MetaMathOctopus on MGSM to validate the generalization of our findings.
Appendix E Layer-wise Changes in the Number of Different Types of Neurons During Alignment
Figure 11 shows the same pattern as the results in the main text, thereby supporting its generalization. Furthermore, for each dataset, we examine the layer-wise changes in the number of different types of neurons in the process of alignment in Figure 1213 and 14. The notation $\text{ckpt-}x$” denotes the model checkpoint obtained after $x$ training steps.
Appendix F Spontaneous Multilingual Alignment
We report the accuracy of the base model on the MGSM and MSVAMP benchmarks in Table 5. The expected accuracy of random guessing is 50.0%. ”X/Y $\Rightarrow$ T” indicates that languages X and Y are used for alignment. The best results for each language are highlighted. It is evident that models trained on multilingual translation data substantially outperform the original model across a wide range of languages, indicating that multilingual training significantly enhances the model’s multilingual capabilities.
Appendix G Limitations
@@
Figure 6: PPL changes of MistralMathOctopus on MGSM when deploying the algorithm proposed by Tang et al. (2024).
@@
(c) Base - Language-specific & -related Neurons
(d) Aligned - Language-specific & -related Neurons
@@
(a) Base - MGSM
(b) Aligned - MGSM
Figure 8: PPL changes of MetaMathOctopus on MGSM after deactivating language-specific neurons and language-related neurons. “Base” indicates the results of the base model. “Aligned” indicates the results of the aligned model.
@@
(a) Before alignment
(b) After alignment
Figure 10: Numbers of neurons shared by NN languages before and after alignment.
@@
Figure 14: Layer-wise changes in the number of general neurons of MistralMathOctopus during the alignment.